# Inferring entropy production in many-body systems using nonequilibrium MaxEnt

### v2

@@ Abstract
We propose a method for inferring entropy production (EP) in high-dimensional stochastic systems, including many-body systems and non-Markovian systems with long memory. Standard techniques for estimating EP become intractable in such systems due to computational and statistical limitations. We infer trajectory-level EP and lower bounds on average EP by exploiting a nonequilibrium analogue of the Maximum Entropy principle, along with convex duality. Our approach uses only samples of trajectory observables (such as spatiotemporal correlation functions). It does not require reconstruction of high-dimensional probability distributions or rate matrices, nor any special assumptions such as discrete states or multipartite dynamics. It may be used to compute a hierarchical decomposition of EP, reflecting contributions from different kinds of interactions, and it has an intuitive physical interpretation as a thermodynamic uncertainty relation. We demonstrate its numerical performance on a disordered nonequilibrium spin model with 1000 spins and a large neural spike-train dataset.
@@
The central quantity of interest in nonequilibrium thermodynamics is entropy production (EP) [1]. EP characterizes departure from thermodynamic equilibrium in microscopic physical systems, where it quantifies the dissipation of thermodynamic free energy. More generally, EP provides an information-theoretic measure of temporal irreversibility, including in meso- and macroscopic systems [2, 3, 4, 5]. There is great interest in quantifying EP in high-dimensional complex systems with many degrees of freedom and/or long non-Markovian memory, such as nonequilibrium disordered networks [6], biological active matter [7, 8], and neural systems [9, 10, 11, 12, 13], but this is computationally and statistically difficult. Standard methods require reconstruction of trajectory probability distributions, an intractable problem for systems with many spatiotemporal degrees of freedom.
In this paper, we propose a method to infer EP in high-dimensional stochastic systems. To do so, we introduce a variational principle that is the nonequilibrium analogue of the well-known Maximum Entropy Principle (MaxEnt) from equilibrium statistical physics. The dual form of this variational principle leads to a tractable optimization problem whose solution provides an approximation of trajectory-level EP and a lower bound on the average EP — directly from trajectory samples and without explicit knowledge of trajectory probabilities. The dual principle also provides a higher-order thermodynamic uncertainty relation (TUR) [14, 15], which we use to derive further simple bounds on EP.
Entropy production.— We consider a nonequilibrium stochastic system measured at discrete time points. The system’s (continuous or discrete) state at any given time is a vector $\bm{x}_{t}=(x_{1,t},x_{2,t},\dots,x_{N,t})$ of dimension $N$. These dimensions may represent physical degrees of freedom, e.g., physical spins or particles in a many-body system. They may also represent population counts (as in a chemical master equation), indicator functions for macrostates, coarse-grained projections, or other types of variables.
The system’s stochastic dynamics over sequences of $T+1$ discrete timesteps are described by a “forward” trajectory distribution $p(\vec{{\bm{x}}})$, where each trajectory $\vec{{\bm{x}}}=({\bm{x}}_{0},{\bm{x}}_{1},\dots,{\bm{x}}_{T})$ is a sequence of $T+1$ states. In general, $T=1$ suffices for stationary first-order-Markovian systems, while larger $T$ may describe non-Markovian and non-stationary systems. The system is also associated with a “reverse” trajectory distribution $\tilde{p}(\vec{{\bm{x}}})$. The trajectory EP of trajectory $\vec{{\bm{x}}}$ and the average EP across all trajectories are defined as [1]
$\sigma(\vec{{\bm{x}}}):=\ln\frac{p(\vec{{\bm{x}}})}{\tilde{p}(\vec{{\bm{x}}})}\,,\qquad\Sigma:=D(p\|\tilde{p})=\left\langle\ln\frac{p(\vec{{\bm{x}}})}{\tilde{p}(\vec{{\bm{x}}})}\right\rangle_{p}\,,$ (1)
where $\langle\ldots\rangle_{p}=\sum_{\vec{{\bm{x}}}}\ldots p(\vec{{\bm{x}}})$ is the expectation over $p$ and $D(\cdot\|\cdot)$ is the Kullback-Leibler (KL) divergence. EP quantifies the violation of time-reversal symmetry, and it vanishes in equilibrium when trajectory statistics are indistinguishable under the forward and reverse distributions.
@@
EP decomposition.— For high-dimensional systems with many degrees of freedom (large $N$) or long memory (large $T$), the set of possible trajectories is exponentially large. In such cases, estimating trajectory probabilities $p(\vec{{\bm{x}}})$ from empirical data becomes statistically infeasible. This prevents estimation of trajectory-level and expected EP by direct evaluation of (1).
We propose a lower bound on EP based on a nonequilibrium generalization of the MaxEnt variational principle. We assume that we have access to $d$ trajectory observables, represented by a vector-valued function $\bm{g}(\vec{{\bm{x}}})\in\mathbb{R}^{d}$. Natural choices of observables include low-order interaction terms, such as 1-point ($x_{i,t}$) and 2-point ($x_{i,t}x_{j,t^{\prime}}$) correlations. We do not require antisymmetric observables $\bm{g}(\vec{{\bm{x}}})=-\bm{g}(\reflectbox{$\vec{\reflectbox{${\bm{x}}$}}$})$.
To quantify the EP captured by the temporal irreversibility of observables $\bm{g}$, we find the distribution that has the smallest KL divergence to the reverse process, subject to having the same observable expectations as the forward process:
${\Sigma}_{\bm{g}}$ $:=\min_{q}D(q\|\tilde{p})\quad\text{subject to}\quad\left\langle\bm{g}\right\rangle_{q}=\left\langle\bm{g}\right\rangle_{p}.$ (2)
${\Sigma}_{\bm{g}}$ is always nonnegative, and it vanishes only when the expectation of $\bm{g}$ is the same under the forward and reverse processes. Furthermore, it is straightforward to show that ${\Sigma}_{\bm{g}}\leq D(p_{G}\|\tilde{p}_{G})\leq{\Sigma}$, where $p_{G}(\bm{g}^{\prime})=\sum_{\vec{{\bm{x}}}}p(\vec{{\bm{x}}})\delta_{\bm{g}^{\prime},\bm{g}(\vec{{\bm{x}}})}$ and $\tilde{p}_{G}(\bm{g}^{\prime})=\sum_{\vec{{\bm{x}}}}\tilde{p}(\vec{{\bm{x}}})\delta_{\bm{g}^{\prime},\bm{g}(\vec{{\bm{x}}})}$ indicate the distributions of $\bm{g}$ under the forward and backward processes (see Supplemental Material (SM) [16]). In general, estimating $D(p_{G}\|\tilde{p}_{G})$ is as difficult as estimating ${\Sigma}$, since it requires inference of the high-dimensional distributions $p_{G}$ and $\tilde{p}_{G}$.
On the other hand, by exploiting a remarkable information-theoretic duality, we may calculate ${\Sigma}_{\bm{g}}$ without explicitly inferring any probability distributions. Specifically, as shown in the End Matter, the optimization problem (2) has the dual formulation
@@
${\Sigma}_{\bm{g}}=\max_{{\bm{\theta}}\in\mathbb{R}^{d}}\left({\bm{\theta}}^{\top}\left\langle\bm{g}\right\rangle_{p}-\ln\langle e^{{\bm{\theta}}^{\top}\bm{g}}\rangle_{\tilde{p}}\right).$ (3)
@@
Eq. (3) is a tractable convex optimization problem over the Lagrangian multipliers ${\bm{\theta}}$ that enforce the $d$ expectation constraints on $\bm{g}$. Importantly, the objective depends only on the expectations of $\bm{g}(\vec{{\bm{x}}})$ under $p$ and $e^{{\bm{\theta}}^{\top}\bm{g}(\vec{{\bm{x}}})}$ under $\tilde{p}$, which can be easily computed from empirical samples of the observables.
@@
As an example, consider a system with 1000 binary spins measured at two timepoints. A naive estimate of ${\Sigma}$ would require inference of $\sim 2^{1000}$ probabilities. Instead, with $\bm{g}$ encoding 2-point correlations, Eq. (3) results in a tractable unconstrained optimization problem over $\sim 1000^{2}$ variables.
@@
So far, we made no assumptions regarding multipartite structure (i.e., only a single variable can change at a time). However, as we show in End Matter, for multipartite dynamics and multipartite observables, Eq. (3) can be split into $N$ separate optimization problems, each one involving only a proportional fraction of the variables and observations. This can dramatically reduce the amount of computation and memory required to numerically solve Eq. (3).
@@
The distribution $q^{*}$ that optimizes (2) lies in the exponential family,
$q_{{\bm{\theta}}}(\vec{{\bm{x}}})=\tilde{p}(\vec{{\bm{x}}})e^{{{\bm{\theta}}}^{\top}\bm{g}(\vec{{\bm{x}}})-\ln\langle e^{{{\bm{\theta}}}^{\top}\bm{g}}\rangle_{\tilde{p}}}.$ (4)
@@
We can interpret the optimal ${\bm{\theta}}^{*}$ in Eq. (3) as the optimal parameters in the maximum likelihood (ML) problem ${\bm{\theta}}^{*}=\operatorname*{arg\,min}_{{\bm{\theta}}}D(p\|q_{{\bm{\theta}}})$, since $q_{{\bm{\theta}}^{*}}=q^{*}$ satisfies the constraints in (2) and achieves the optimum objective value (see End Matter). This interpretation also leads to the information-geometric Pythagorean theorem [17], which decomposes EP into two nonnegative KL terms:
@@
The term ${\Sigma}^{\perp}_{\bm{g}}:={\Sigma}-{\Sigma}_{\bm{g}}=D(p\|q^{*})$ quantifies the EP not captured by ${\Sigma}_{\bm{g}}$. It can be expressed in terms of a ML problem, ${\Sigma}^{\perp}_{\bm{g}}=\min_{{\bm{\theta}}\in\mathbb{R}^{d}}D(p\|q_{\bm{\theta}})$. To interpret this expression, we define a larger exponential family $\omega_{f}(\vec{{\bm{x}}}):=\tilde{p}(\vec{{\bm{x}}})e^{f(\vec{{\bm{x}}})}/\langle e^{f}\rangle_{\tilde{p}}$ that contains the forward trajectory distribution $p(\vec{{\bm{x}}})=\omega_{\sigma}(\vec{{\bm{x}}})$. The ML problem can be written as
${\Sigma}^{\perp}_{\bm{g}}=\min_{{\bm{\theta}}\in\mathbb{R}^{d}}D(\omega_{\sigma}\|\omega_{{\bm{\theta}}^{\top}\bm{g}}).$ (6)
From this expression, it is clear that ${\Sigma}^{\perp}_{\bm{g}}=0$ (and ${\Sigma}_{\bm{g}}={\Sigma}$) if and only if trajectory EP is a linear combination of the trajectory observables, i.e., $\sigma(\vec{{\bm{x}}})={\bm{\theta}}^{*\top}\bm{g}(\vec{{\bm{x}}})$ for all $\vec{{\bm{x}}}$. When ${\Sigma}^{\perp}_{\bm{g}}>0$, ${\bm{\theta}}^{*\top}\bm{g}(\vec{{\bm{x}}})$ provides the best possible approximation of $\sigma(\vec{{\bm{x}}})$ in the span of $\bm{g}$.
@@
Eq. (4) implies that the conditional distribution of the optimal $q^{*}$ obeys $q^{*}_{\vec{X}|G}(\vec{{\bm{x}}}|\bm{g}^{\prime})=\tilde{p}_{\vec{X}|G}(\vec{{\bm{x}}}|\bm{g}^{\prime})$, with conditional probabilities given by Bayes’ rule, $p_{\vec{X}|G}(\vec{{\bm{x}}}|\bm{g}^{\prime})=p(\vec{{\bm{x}}})\delta_{\bm{g}^{\prime},\bm{g}(\vec{{\bm{x}}})}/p_{G}(\bm{g}^{\prime})$. Using the chain rule of KL divergence, we may further decompose ${\Sigma}^{\perp}_{\bm{g}}$ into
${\Sigma}^{\perp}_{\bm{g}}=D(p_{G}\|q^{*}_{G})+D(p_{\vec{X}|G}\|\tilde{p}_{\vec{X}|G})\,.$ (7)
@@
The first term in Eq. (7) vanishes when $q^{*}_{G}=p_{G}$; it is the additional EP that could be inferred by constraining all higher-order statistics (not just the mean) in Eq. (2). The second term is the EP that cannot be inferred from any statistics of $\bm{g}$.
Finally, our method supports a hierarchical decomposition of EP, where contributions from interactions of increasing order (e.g., pairwise, triplets) are successively added, resulting in a structure $0\leq{\Sigma}_{1}\leq{\Sigma}_{2}\leq\dots\leq{\Sigma}$, with $\Sigma_{k}=D(q^{*}_{k}\|\tilde{p})$ capturing contributions from functions $\bm{g}(\vec{{\bm{x}}})$ containing interactions up to order $k$. This structure mirrors information-geometric decompositions in equilibrium MaxEnt [17, 18] and allows EP to be broken down into interpretable components, each satisfying $\Sigma_{k}-\Sigma_{k-1}=D(q^{*}_{k}\|q^{*}_{k-1})$.
@@
Thermodynamic uncertainty relations (TURs).— Eq. (3) can be interpreted as a TUR that relates EP and fluctuations of trajectory observables [19]. Consider any trajectory observable that can be expressed as a linear combination $o(\vec{{\bm{x}}})={\bm{\theta}}^{\top}\bm{g}(\vec{{\bm{x}}})$ for some ${\bm{\theta}}\in\mathbb{R}^{d}$. The first term in (3) is the expectation of $o(\vec{{\bm{x}}})$, while the second term is the cumulant generating function (CGF) of $o(\vec{{\bm{x}}})$ under $\tilde{p}$. Thus, ${\Sigma}_{\bm{g}}$ bounds the “fluctuation discounted” expectation of all such $o(\vec{{\bm{x}}})$, giving a higher-order TUR [14] that constrains all cumulants of trajectory observables. This contrasts with quadratic TURs, which constrain only the mean and variance [20, 21, 22].
@@
To relate our approach to quadratic TURs, we expand the CGF as $\ln\langle e^{{\bm{\theta}}^{\top}\bm{g}}\rangle_{\tilde{p}}\approx{\bm{\theta}}^{\top}\left\langle\bm{g}\right\rangle_{\tilde{p}}+{\bm{\theta}}^{\top}\mathsf{K}_{\tilde{p}}{\bm{\theta}}/2$, where $\mathsf{K}_{\tilde{p}}$ is the covariance matrix of $\bm{g}$ under $\tilde{p}$. Plugging into (3) gives the approximate optimizer $\hat{{\bm{\theta}}}$ as the solution to the linear system $\mathsf{K}_{\tilde{p}}\hat{{\bm{\theta}}}=\left\langle\bm{g}\right\rangle_{p}-\left\langle\bm{g}\right\rangle_{\tilde{p}}\equiv\left\langle\bm{g}\right\rangle_{p-\tilde{p}}$. This gives the weaker bound
@@
${\Sigma}_{\bm{g}}\geq\widehat{{\Sigma}}_{\bm{g}}$ $:=\left\langle\bm{g}\right\rangle_{p-\tilde{p}}^{\top}\mathsf{K}_{\tilde{p}}^{-1}\left\langle\bm{g}\right\rangle_{p}-\ln\left\langle e^{\left\langle\bm{g}\right\rangle_{p-\tilde{p}}^{\top}\mathsf{K}_{\tilde{p}}^{-1}\bm{g}}\right\rangle_{\tilde{p}}.$ (8)
@@
$\widehat{{\Sigma}}_{\bm{g}}$ can be shown to correspond to a single Newton-Raphson step for maximizing (3) starting from ${\bm{\theta}}=\bm{0}$ [16]; it has the advantage of not requiring optimization, only the solution a linear system. The bound ${\Sigma}_{\bm{g}}\geq\widehat{{\Sigma}}_{\bm{g}}$ becomes tight when $\bm{g}$ has Gaussian statistics.
We may also compare our results to a quadratic TUR. For stationary systems without odd variables, we find the inequality (see End Matter)
@@
${\Sigma}\geq{\Sigma}_{\bm{g}}^{\textrm{TUR}}:=\frac{1}{2}(\langle\bm{g}\rangle_{p}-\langle\bm{g}\rangle_{\tilde{p}})^{\top}\mathsf{K}_{(p+\tilde{p})/2}^{-1}(\langle\bm{g}\rangle_{p}-\langle\bm{g}\rangle_{\tilde{p}})$ (9)
where $\mathsf{K}_{(p+\tilde{p})/2}$ is the covariance matrix of $\bm{g}$ under the mixture distribution $(p+\tilde{p})/2$. For antisymmetric observables, Eq. (9) may be considered the discrete-time version of the multidimensional TUR from Ref. [22].

### v3

@@ Abstract
We propose a method for inferring entropy production (EP) in high-dimensional stochastic systems, including many-body systems and non-Markovian systems with long memory. Standard techniques for estimating EP become intractable in such systems due to computational and statistical limitations. We infer trajectory-level EP and lower bounds on average EP by exploiting a nonequilibrium analogue of the Maximum Entropy principle, along with convex duality. Our approach uses only samples of trajectory observables, such as spatiotemporal correlations. It does not require reconstruction of high-dimensional probability distributions or rate matrices, nor impose any special assumptions such as discrete states or multipartite dynamics. In addition, it may be used to compute a hierarchical decomposition of EP, reflecting contributions from different interaction orders, and it has an intuitive physical interpretation as a “thermodynamic uncertainty relation.” We demonstrate its numerical performance on a disordered nonequilibrium spin model with 1000 spins and a large neural spike-train dataset.
@@
The central quantity of interest in nonequilibrium thermodynamics is entropy production (EP) [1]. In microscopic physical systems, EP characterizes departure from thermodynamic equilibrium and quantifies the dissipation of thermodynamic free energy. More generally, EP provides an information-theoretic measure of temporal irreversibility, including in meso- and macroscopic systems [2, 3, 4, 5].
Recently, there has been growing interest in thermodynamic inference [6], that is, the problem of estimating EP from empirical measurements of a stochastic system. Various methods have been developed for inferring EP from partial [7, 8, 9, 10, 11, 12, 13, 14] information. For instance, the celebrated thermodynamic uncertainty relation (TUR) [15, 16, 17, 18] bounds EP in terms of the mean and variance of a single current. Other techniques relate EP to the statistics of waiting times [19, 20, 21, 22], observed transitions [23, 24, 11, 12], and counting observables [25]. Many of these methods are designed to infer underlying dissipation from a small number of coarse-grained observables.
In this Letter, we consider thermodynamic inference in the high-dimensional setting, where a large number of observables are available. This is particularly relevant for multivariate measurements of complex systems — such as nonequilibrium disordered networks [26], biological active matter [27, 28], and neural systems [29, 30, 31, 32, 33] — that have many degrees of freedom and/or long non-Markovian memory. A naïve approach requires the estimation of trajectory probability distributions, but in high dimensions, this is usually statistically and numerically infeasible.
To address this challenge, we propose to infer EP using an information-theoretic variational principle, which can be understood as the nonequilibrium analogue of the Maximum Entropy Principle (MaxEnt) in statistical physics. Our variational principle has a simple physical interpretation: it quantifies the irreversibility captured by the expectations of some (possibly large) number of trajectory observables. Importantly, this variational principle has a dual form that leads to a tractable convex optimization problem. This optimization problem gives a lower bound on EP and an estimate of the fluctuating trajectory-level EP — directly from trajectory samples and without explicit use of trajectory probabilities.
As we show, our bound can also be understood as a higher-order thermodynamic uncertainty relation (TUR) [34, 35], which we use to derive further simple bounds on EP. Finally, we show that in multipartite systems, our variational principle can be split into smaller subproblems, greatly improving the performance scaling of our method with system size.
@@
The trajectory EP of trajectory $\bm{x}$}} and the average EP across all trajectories are defined in terms of symmetry breaking between forward and reverse distributions [1],
@@
$\sigma({{\bm{x}}})\vcentcolon=\ln\frac{p({{\bm{x}}})}{\tilde{p}({{\bm{x}}})}\,,\qquad\Sigma\vcentcolon=D(p\|\tilde{p})=\left\langle\ln\frac{p({{\bm{x}}})}{\tilde{p}({{\bm{x}}})}\right\rangle_{p}\,$, (1)
where $\langle\ldots\rangle_{p}=\sum_{{{\bm{x}}}}\ldots p({{\bm{x}}})$ is the expectation over $p$ and $D(\cdot\|\cdot)$ is the Kullback-Leibler (KL) divergence. EP vanishes in equilibrium, when trajectory statistics are indistinguishable under the forward and reverse distributions.
[…]
Our goal is to estimate $\sigma({{\bm{x}}})$ and $\Sigma$ from forward and reverse samples of observables $\bm{g}$. The overall irreversibility encoded in these observables is quantified by the KL divergence between the distributions of $\bm{g}$ under the forward and reverse processes, ${\Sigma}_{\bm{g}}^{\mathrm{DPI}}\vcentcolon=D(p_{G}\|\tilde{p}_{G})$, where $p_{G}(\bm{g}^{\prime})=\left\langle\delta_{\bm{g}^{\prime},\bm{g}({{\bm{x}}})}\right\rangle_{p}$ and $\tilde{p}_{G}(\bm{g}^{\prime})=\left\langle\delta_{\bm{g}^{\prime},\bm{g}({{\bm{x}}})}\right\rangle_{\tilde{p}}$. The data processing inequality (DPI) implies that ${\Sigma}_{\bm{g}}^{\mathrm{DPI}}\leq{\Sigma}$, with equality when $\bm{g}({{\bm{x}}})$ is an invertible function [36].
Estimating ${\Sigma}_{\bm{g}}^{\mathrm{DPI}}$ is very difficult for high-dimensional observables $\bm{g}$. Instead, we bound EP using a nonequilibrium generalization of the MaxEnt variational principle. Specifically, we choose the distribution that minimizes KL divergence to the reverse process while matching the forward-process expectations of the observables:
${\Sigma}_{\bm{g}}$ $\vcentcolon=\min_{q}D(q\|\tilde{p})\quad\text{subject to}\quad\left\langle\bm{g}\right\rangle_{q}=\left\langle\bm{g}\right\rangle_{p}.$ (2)
This quantity obeys the bounds $0\leq{\Sigma}_{\bm{g}}\leq{\Sigma}_{\bm{g}}^{\mathrm{DPI}}$, as shown in the Supplemental Material (SM) [37]. ${\Sigma}_{\bm{g}}$ quantifies the irreversibility captured by observable expectations, and it vanishes if and only if the expectation of $\bm{g}$ is the same under the forward and reverse processes. Large-deviations theory gives ${\Sigma}_{\bm{g}}$ a physical interpretation in terms of fluctuations: given $n$ sample trajectories from the reverse process, the probability that the empirical average of observables $\bm{g}$ is equal to $\left\langle\bm{g}\right\rangle_{p}$ scales as $\asymp e^{-n{\Sigma}_{\bm{g}}}$ [38].
Importantly, by exploiting a remarkable information-theoretic duality, we may calculate ${\Sigma}_{\bm{g}}$ without optimizing or inferring any probability distributions. As shown in SM [37], the optimization problem (2) has the dual formulation
[…]
@@
Eq. (3) is an unconstrained convex optimization problem over Lagrangian multipliers $\bm{\theta}$} that enforce the $d$ expectation constraints on $\bm{g}$. The objective depends only on the expectations of $\bm{g}$ and $e^{{\bm{\theta}}^{\top}\bm{g}}$ under the forward and backward processes, which can be easily computed from empirical samples of $\bm{g}$. In the special case of antisymmetric observables and a steady-state system without odd variables, Eq. (3) can be estimated using only forward samples since $\ln\langle e^{{\bm{\theta}}^{\top}\bm{g}}\rangle_{\tilde{p}}=\ln\langle e^{-{\bm{\theta}}^{\top}\bm{g}}\rangle_{p}$. As we discuss in more detail below, a related (but different) variational bound on EP was proposed in Refs. [39, 40].
@@
As an example, consider a system with 1000 binary spins measured at two timepoints, with $\bm{g}$ encoding two-point correlations $x_{i,t}x_{j,t^{\prime}}$. Eq. (3) is an unconstrained convex optimization problem over $1000^{2}$ variables, which can be solved using standard numerical techniques. Conversely, a naïve estimate of $\Sigma$} or ${\Sigma}_{\bm{g}}^{\mathrm{DPI}}$ requires inferrence of $2^{1000}$ probabilities.
@@
So far, we have made no assumptions regarding multipartite structure. As we show in End Matter, in the presence of multipartite structure, where only a single observable can be active in a given trajectory, Eq. (3) can be split into smaller optimization problems, each involving a proportional fraction of variables and samples. This dramatically reduces the computation and memory required to solve Eq. (3) (see End Matter for an example).
@@
Maximum likelihood EP decomposition.— As we show in the SM [37], the distribution $q^{*}$ that optimizes (2) belongs to an exponential family,
$q_{{\bm{\theta}}}({{\bm{x}}})=\tilde{p}({{\bm{x}}})e^{{{\bm{\theta}}}^{\top}\bm{g}({{\bm{x}}})-\ln\langle e^{{{\bm{\theta}}}^{\top}\bm{g}}\rangle_{\tilde{p}}}$. (4)
Specifically, $q^{*}=q_{{\bm{\theta}}^{*}}$ for optimal parameters $\bm{\theta}$$^{*}$ that assign maximum likelihood (ML) to forward trajectories:
${\bm{\theta}}^{*}=\operatorname*{arg\,max}_{{\bm{\theta}}\in\mathbb{R}^{d}}\,\big{\langle}\ln q_{{\bm{\theta}}}({{\bm{x}}})\big{\rangle}_{p}=\operatorname*{arg\,min}_{{\bm{\theta}}\in\mathbb{R}^{d}}\,D(p\|q_{{\bm{\theta}}}).$ (5)
@@
This also leads to the information-geometric Pythagorean theorem [41], which decomposes EP into two nonnegative terms:
@@
${\Sigma}^{\perp}_{\bm{g}}\vcentcolon={\Sigma}-{\Sigma}_{\bm{g}}=D(p\|q^{*})$ quantifies the EP not captured by ${\Sigma}_{\bm{g}}$, expressed via the ML problem ${\Sigma}^{\perp}_{\bm{g}}=\min_{{\bm{\theta}}}D(p\|q_{\bm{\theta}})$.
@@
We may further decompose ${\Sigma}^{\perp}_{\bm{g}}$ by using that Eq. (4) implies that $q^{*}$ obeys $q^{*}_{X|G}({{\bm{x}}}|\bm{g}^{\prime})=\tilde{p}_{X|G}({{\bm{x}}}|\bm{g}^{\prime})$, where conditional probabilities are computed using Bayes’ rule, $\tilde{p}_{X|G}({{\bm{x}}}|\bm{g}^{\prime})=\tilde{p}({{\bm{x}}})\delta_{\bm{g}^{\prime},\bm{g}({{\bm{x}}})}/\tilde{p}_{G}(\bm{g}^{\prime})$. Using the chain rule of KL divergence, we may write ${\Sigma}^{\perp}_{\bm{g}}$ as a sum of two terms:
${\Sigma}^{\perp}_{\bm{g}}=\underbrace{D(p_{G}\|q^{*}_{G})}_{{\Sigma}_{\bm{g}}^{\mathrm{DPI}}-{\Sigma}_{\bm{g}}}+\underbrace{D(p_{X|G}\|\tilde{p}_{X|G})}_{{\Sigma}-{\Sigma}_{\bm{g}}^{\mathrm{DPI}}}\,.$ (7)
@@
The first term is the additional EP that could be inferred by constraining all statistics (not just the mean) of $\bm{g}$ in Eq. (2). The second term is the EP that cannot be inferred from any statistics of $\bm{g}$.
[…]
Finally, our method gives an estimate of trajectory EP,
$\sigma_{{\bm{\theta}}^{*}}({{\bm{x}}})\vcentcolon=\ln\frac{q_{{\bm{\theta}}^{*}}({{\bm{x}}})}{\tilde{p}({{\bm{x}}})}={\bm{\theta}}^{*\top}\bm{g}({{\bm{x}}})-\ln\langle e^{{\bm{\theta}}^{*\top}\bm{g}}\rangle_{\tilde{p}}\,$. (8)
This estimator satisfies $\langle\sigma_{{\bm{\theta}}^{*}}\rangle_{p}={\Sigma}_{\bm{g}}$, and it may be shown to provide the best approximation of trajectory EP $\sigma({{\bm{x}}})$ within the exponential family $q_{\bm{\theta}}$. Specifically, if we use $\sigma_{{\bm{\theta}}}({{\bm{x}}})=\ln[q_{{\bm{\theta}}}({{\bm{x}}})/{\tilde{p}}({{\bm{x}}})]$ to indicate the approximation given by parameters $\bm{\theta}$} and $\delta\sigma_{{\bm{\theta}}}({{\bm{x}}})=\sigma({{\bm{x}}})-\sigma_{{\bm{\theta}}}({{\bm{x}}})$ the corresponding residual, then $\sigma_{{\bm{\theta}}^{*}}$ minimizes the expected error  [^1] Eq. (9) is derived by using $\delta\sigma_{{\bm{\theta}}}({{\bm{x}}})=\ln p({{\bm{x}}})/q_{{\bm{\theta}}}({{\bm{x}}})$ and $\langle e^{-\delta\sigma_{{\bm{\theta}}}}\rangle_{p}=1$.:
$D(p\|q_{{\bm{\theta}}})=\langle\delta\sigma_{{\bm{\theta}}}+e^{-\delta\sigma_{{\bm{\theta}}}}-1\rangle_{p}.$ (9)
Here, $\delta\sigma_{{\bm{\theta}}}+e^{-\delta\sigma_{{\bm{\theta}}}}-1\geq 0$ is an information-theoretic loss function that converges to squared error $(\delta\sigma_{{\bm{\theta}}})^{2}/2$ as $\delta\sigma_{{\bm{\theta}}}\to 0$.
[…]
$\sigma({{\bm{x}}})=\sigma_{{\bm{\theta}}^{*}}({{\bm{x}}})={\bm{\theta}}^{*\top}[\bm{g}({{\bm{x}}})-\bm{g}(\tilde{\bm{x}})]/2\,$, (10)
which simplifies further to ${\bm{\theta}}^{*\top}\bm{g}({{\bm{x}}})$ when $\bm{g}$ is antisymmetric.
@@
Thermodynamic uncertainty relations (TURs).— Eq. (3) can be interpreted as a TUR that relates EP and fluctuations of trajectory observables [44]. Consider any trajectory observable that can be expressed as a linear combination $o({{\bm{x}}})={\bm{\theta}}^{\top}\bm{g}({{\bm{x}}})$ for some ${\bm{\theta}}\in\mathbb{R}^{d}$. The first term in (3) is the expectation of $o({{\bm{x}}})$, while the second term is the cumulant generating function (CGF) of $o({{\bm{x}}})$ under $\tilde{p}$. Thus, ${\Sigma}_{\bm{g}}$ bounds the fluctuation-discounted expectation of all such $o({{\bm{x}}})$, and it can be understood as a higher-order TUR [34] that constrains all cumulants of trajectory observables. This contrasts with quadratic TURs, which constrain only the mean and variance [45, 18, 46].
@@
To relate our approach to quadratic TURs, we expand the CGF as $\ln\langle e^{{\bm{\theta}}^{\top}\bm{g}}\rangle_{\tilde{p}}\approx{\bm{\theta}}^{\top}\left\langle\bm{g}\right\rangle_{\tilde{p}}+{\bm{\theta}}^{\top}\mathsf{K}_{\tilde{p}}\,{\bm{\theta}}/2$, where $\mathsf{K}_{\tilde{p}}$ is the covariance matrix of $\bm{g}$ under $\tilde{p}$. Plugging into (3) gives the approximate optimizer $\hat{{\bm{\theta}}}$ as the solution to the linear system $\mathsf{K}_{\tilde{p}}\,\hat{{\bm{\theta}}}=\left\langle\bm{g}\right\rangle_{p}-\left\langle\bm{g}\right\rangle_{\tilde{p}}\equiv\left\langle\bm{g}\right\rangle_{p-\tilde{p}}$, giving the weaker bound
@@
${\Sigma}_{\bm{g}}\geq\widehat{{\Sigma}}_{\bm{g}}$ $\vcentcolon=\left\langle\bm{g}\right\rangle_{p-\tilde{p}}^{\top}\mathsf{K}_{\tilde{p}}^{-1}\left\langle\bm{g}\right\rangle_{p}-\ln\left\langle e^{\left\langle\bm{g}\right\rangle_{p-\tilde{p}}^{\top}\mathsf{K}_{\tilde{p}}^{-1}\bm{g}}\right\rangle_{\tilde{p}}$. (11)
@@
$\widehat{{\Sigma}}_{\bm{g}}$ can be shown to correspond to a single Newton-Raphson step for maximizing (3) starting from ${\bm{\theta}}=\bm{0}$ [37]. This bound has the advantage of not requiring optimization, only the solution of a linear system. The inequality ${\Sigma}_{\bm{g}}\geq\widehat{{\Sigma}}_{\bm{g}}$ is tight when $\bm{g}$ has Gaussian statistics.
We may also compare our results to existing TURs. We consider the special case of steady-state system without odd variables and antisymmetric $\bm{g}$. In this case, EP may be bounded as (see End Matter)
@@
${\Sigma}\geq{\Sigma}_{\bm{g}}^{\textrm{TUR}}\vcentcolon=\ln\left(1+2\langle\bm{g}\rangle_{p}^{\top}\mathsf{K}_{p}^{-1}\langle\bm{g}\rangle_{p}\right)$ (12)
This bound approaches the usual quadratic form, ${\Sigma}_{\bm{g}}^{\textrm{TUR}}\simeq 2\langle\bm{g}\rangle_{p}^{\top}\mathsf{K}_{p}^{-1}\langle\bm{g}\rangle_{p}$ for sufficiently small $\langle\bm{g}\rangle_{p}^{\top}\mathsf{K}_{p}^{-1}\langle\bm{g}\rangle_{p}$. This regime holds in the short-time limit, since $\langle\bm{g}\rangle_{p}$ scales with the observation period.